OpenStudy (anonymous):

Anyone take 12th grade PreCalc @ Connections?

RhondaSommer (rhondasommer):

Im not in 12th grade. nor do i use connections, but i am really good at precalc. would you like my help?

OpenStudy (anonymous):

yes PLEASE! we can do different questions on different posts so I can give you medals!

OpenStudy (anonymous):

@RhondaSommer find the value of csc theta for angle theta in standard position if the point at (3, -1) lies on its terminal side.

OpenStudy (anonymous):

answer choices: -sqrt10 -sqrt10/10 sqrt10/3 -3

OpenStudy (anonymous):

thank you so much for even trying you are a lifesaver!!

RhondaSommer (rhondasommer):

yeah :)

RhondaSommer (rhondasommer):

id say a though

RhondaSommer (rhondasommer):

\[3^2+1^2=h^2\]\[9+1=h^2\]\[10=h^2\]\[h=sqrt10\]

OpenStudy (anonymous):

is that negative? because I don't have that answer

RhondaSommer (rhondasommer):

answer choices: -sqrt10 -sqrt10/10 sqrt10/3 -3

RhondaSommer (rhondasommer):

sqrt10 is the answer

OpenStudy (anonymous):

so the first one? its negative.

RhondaSommer (rhondasommer):

the first one is a negative?

OpenStudy (anonymous):

yes

RhondaSommer (rhondasommer):

i mean i could be wrong.. f point (3, -1) lies on the terminal side of an angle A in standard position, then i found the answe and its sqrt 10

OpenStudy (anonymous):

the first answer choice, I mean.

RhondaSommer (rhondasommer):

go for it

OpenStudy (anonymous):

okay. another maybe?

RhondaSommer (rhondasommer):

sure

RhondaSommer (rhondasommer):

tag me